Napoleon’s cannon, having a muzzle velocity of 126 m/s, is fired at a 0 degree angle off a cliff. The ball lands 739 m away from the cannon. How far above ground level is the cannon?
1
Expert's answer
2020-09-28T06:04:11-0400

Let us write equations of motion of cannon:

"y(t) = H - \frac{g t^2}{2}", "x(t) = v_0 t" , where "H" is the height of the cliff, that we need to find, "v_0 = 126 \frac{m}{s}" is the muzzle velocity, "S = 739 m" is the horizontal distance, that the ball covered.

When the ball reaches the ground in time "t_1", "y(t_1) = 0", from where using equation for y:"0 = H - \frac{g t_1^2}{2}" , hence the height of the cliff is "H = \frac{g t_1^2}{2}".

Let us find time "t_1" using equation for x:

At time "t_1", "x(t_1) = S = v_0 t_1", from where "t_1 = \frac{S}{v_0}".

Substituting last expression for "t_1" into formula for "H", obtain:

"H = \frac{g t_1^2}{2} = \frac{g}{2} \frac{S^2}{v_0^2} \approx 168.7 m"